how much water must be evaporated from a 300 L tank of a 2% salt solution to obtain a 5% solution?
Answer by Paul(988)

Let the solution be x
2(300)=5(300-x)
600=1500-5x
-5x=-900
x=180

Hence, about 180L of solution must be evaporated.
PAul.